OpenStudy (anonymous):

Help Please !! Choose the equivalent system of linear equations that will produce the same solution as the one given below. 4x - 2y = 6 2x + y = 5 -4x - 2y = 10 -4y = 4 -4x - 5y = -1 -7y = 5 3x + 2y = 6 7x = 12 4x + 2y = 10 8x = 16

OpenStudy (anonymous):

anyobdy there ??

OpenStudy (phi):

This looks like it takes time. I would solve the system they "gave" as a first step. can you do that ?

OpenStudy (anonymous):

i dont know how to @phi

OpenStudy (phi):

4x - 2y = 6 2x + y = 5 multiply the second equation by 2. (that means *all* terms, on both sides by 2) can you do that ?

OpenStudy (anonymous):

which would be 4x + 2x = 5

OpenStudy (phi):

almost. that second term should be 2y (right?) and when we double the left side we have to double the right side to keep it an equation in other words, don't forget the 5 try again.

OpenStudy (anonymous):

ok so 4x + 2y = 10

OpenStudy (phi):

4x - 2y = 6 4x + 2y = 10 now add the two equations. which just means write them as 4x+4x -2y +2y = 6+16 can you simplify that ?

OpenStudy (anonymous):

hold on let me try

OpenStudy (phi):

**4x+4x -2y +2y = 6+10

OpenStudy (anonymous):

8x - y = 16

OpenStudy (phi):

ok, except -2y + 2y is not -y 2 y's take away 2 y's is how many y's ?

OpenStudy (anonymous):

it would be 0

OpenStudy (phi):

yes, and 4x+4x -2y +2y = 6+10 simplifies to 8x= 16 now divide both sides by 8

OpenStudy (anonymous):

2 ?

OpenStudy (phi):

you should write both sides 8x/8 = 16/8 on the left side 8/8 * x is 1*x = x on the right side 16/8 is 2 we get x=2

OpenStudy (anonymous):

so it would be D ?

OpenStudy (phi):

this is a painful question. so far we know x=2 use one of the two equations they gave us example 2x + y = 5 (I picked this because it has smaller numbers, but the other works also) replace x with 2 2*2 + y =5 can you "solve" for y? I would simplify 2*2 to 4 then add -4 to both sides

OpenStudy (anonymous):

Dang.. i dont know how to do all that i figured it would be B since we got 8x= 16

OpenStudy (phi):

you do it like lots of things. 1 step at a time 2x + y = 5 we figured out x=2 replace x with 2: 2*2 + y = 5 what is 2*2 ?

OpenStudy (anonymous):

4

OpenStudy (phi):

ok. 2*2 + y = 5 becomes 4+ y = 5 to "get rid" of the 4, add -4 (to both sides to keep it an equation) 4-4 + y = 5 - 4 simplify

OpenStudy (anonymous):

y=1

OpenStudy (phi):

ok, the "solution" to the given system is x=2, y=1 we need to try those numbers in the choices and see if they work for one of the choices (they better work for one of them) Let's test the first choice: -4x - 2y = 10 -4y = 4 the second equation is easier to test. if we replace y with 1 we get -4*1 =? -4 -4 =? 4 .... Nope , that is not right. First choice is not the answer.

OpenStudy (anonymous):

Ohh its b

OpenStudy (phi):

can you test -4x - 5y = -1 -7y = 5 test the second equation with y=1. does it work?

OpenStudy (anonymous):

yes

OpenStudy (phi):

can you show how you did it?

OpenStudy (anonymous):

ohh no im not saying its b it was a typo i was trying to say a -4y=4

OpenStudy (phi):

yes, but when you try y=1 you get -4*1 =4 and -4*1 is -4 and -4 does not equal 4 It does not work.

OpenStudy (anonymous):

omg im geting frustrated w/ this

OpenStudy (phi):

It's a painful problem. But it's doable. Do you follow how we tested -4y= 4 to see if it's true if y is 1 ? we found it did NOT work.

OpenStudy (anonymous):

yes sp im guessing its 7x=12 (c)

OpenStudy (anonymous):

so*

OpenStudy (phi):

no need to guess. remember x=2 is the solution for our system does x=2 work for 7x=12 ?

OpenStudy (anonymous):

ugh no it doesnt :(

OpenStudy (phi):

that's good. we can cross of that choice.

OpenStudy (anonymous):

okay and we can also cross out a right ?

OpenStudy (phi):

test the last one 4x + 2y = 10 8x = 16 use the last equation . Does it work for x=2 ?

OpenStudy (anonymous):

yes thats what i said the first time it was D

OpenStudy (phi):

ok. but the first equation has to also work. 4x+2y must equal 10 when we use x=2 and y=1 does it ?

OpenStudy (anonymous):

no

OpenStudy (phi):

4*2 +2*1 first do the multiplies then add

OpenStudy (anonymous):

8 + 3 = 11

OpenStudy (phi):

2*1 is different from 2+1

OpenStudy (anonymous):

ohh im sorry it equals 10

OpenStudy (phi):

yes, so choice D is the answer. A tough question (lots of work)

OpenStudy (anonymous):

omg thank you so much !

OpenStudy (phi):

yw